OpenStudy (anonymous):

How many gallons of a 90%solution must be mixed with 90 gallons of 30% antifreeze to get a mixture that is 80% antifreeze?

OpenStudy (anonymous):

I got 432 is that correct?

OpenStudy (anonymous):

lets call the amount of 90% solution \(x\) so it will contain \(.9x\) antifreeze you have 90 gallons of 30% solution, so it contains \(.3\times 90=27\) gallons of antifreeze when you add \(x\) gallons to the 90 gallons you will have \(x+90\) gallons that you want to be 80% antifreeze, so set \[.9x+27=.8(90+x\] and solve for \(x\)

OpenStudy (anonymous):

that is not what i get, no

OpenStudy (anonymous):

\[.9x+27=.8(90+x)\] get rid of annoying decimal by multplying by 10 and get \[9x+270=8(90+x)\]

OpenStudy (anonymous):

ugh Im lost ive done this prob a million times

OpenStudy (anonymous):

\[9x+270=720+8x\] \[x+270=720\] \[x=450\] is what i get

OpenStudy (anonymous):

yeah they are a pain any question about any steps?

OpenStudy (anonymous):

No I think i WAS JUST MISCALCULATING